# Analysis of mean field games via Fokker-Planck-Kolmogorov equations:

### v1

@@ 3. Auxiliary results
$\int_{\mathbb{R}^{d}}\psi(y)\,\sigma_{t}(dy)-\int_{\mathbb{R}^{d}}\psi(y)\,\nu(dy)=\\ \int_{0}^{t}\int_{\mathbb{R}^{d}}L_{\mu,0}\psi(y,s)\,\sigma_{s}(dy)\,ds+\int_{0}^{t}\int_{\mathbb{R}^{d}\times U}\langle Q(y,s,\mu)v,\nabla\psi(y)\rangle\Pi_{s}(dydv)\,ds$
holds for every $\psi\in C_{0}^{\infty}(\mathbb{R}^{d})$ and all $t\in[0,T]$. Denote by $\Pi_{y,s}(dv)$ the conditional measures for $\Pi_{s}(dydv)$ with respect to $\sigma_{s}$. There exists a Borel function $(y,s)\mapsto v(y,s)$ such that the equality
@@
$\int_{0}^{t}\int_{\mathbb{R}^{d}\times U}\langle Q(y,s,\mu)v,\nabla\psi(y)\rangle\Pi_{s}(dydv)\,ds=\int_{0}^{t}\int_{\mathbb{R}^{d}}\langle Q(y,t,\mu)v(y,s),\nabla\psi(y)\rangle\sigma_{s}(dy)\,ds,$
@@
$\mathbb{E}\Bigl{[}\int_{0}^{T}f(v_{t},y_{t},t,\mu)\,dt+g(y_{T},\mu)\Bigr{]}=\int_{0}^{T}\int_{\mathbb{R}^{d}\times U}f(v,y,t,\mu)\Pi_{t}(dydv)\,dt+\int_{\mathbb{R}^{d}}g(y,\mu)\sigma_{T}(dy).$
@@
$\int_{0}^{T}\int_{\mathbb{R}^{d}\times U}f(v,y,t,\mu)\Pi_{t}(dydv)\,dt\geq\int_{0}^{T}\int_{\mathbb{R}^{d}\times U}f(v(y,t),y,t,\mu)\sigma_{t}(dy)\,dt.$
Now we can apply Theorem 2.1. ∎
@@ Acknowledgements.
The authors are grateful to Prof. Vladimir I. Bogachev, Prof. Yurii V. Averboukh and Tikhon I. Krasovitskii for fruitful discussions and valuable remarks.

### v2

@@ 3. Auxiliary results
$\int_{\mathbb{R}^{d}}\psi(y)\,\sigma_{t}(dy)-\int_{\mathbb{R}^{d}}\psi(y)\,\nu(dy)=\\ \int_{0}^{t}\int_{\mathbb{R}^{d}}L_{\mu,0}\psi(y,s)\,\sigma_{s}(dy)\,ds+\int_{U\times\mathbb{R}^{d}\times[0,t]}\langle Q(y,s,\mu)v,\nabla\psi(y)\rangle\,\Pi(dvdyds).$
holds for every $\psi\in C_{0}^{\infty}(\mathbb{R}^{d})$ and all $t\in[0,T]$. Denote by $\Pi_{y,s}(dv)$ the conditional measures for $\Pi$ with respect to $\sigma_{s}\,ds$. There exists a Borel function $(y,s)\mapsto v(y,s)$ such that the equality
@@
$\int_{U\times\mathbb{R}^{d}\times[0,t]}\langle Q(y,s,\mu)v,\nabla\psi(y)\rangle\,\Pi(dvdyds)=\int_{0}^{t}\int_{\mathbb{R}^{d}}\langle Q(y,t,\mu)v(y,s),\nabla\psi(y)\rangle\sigma_{s}(dy)\,ds,$
@@
$\mathbb{E}\Bigl[\int_{0}^{T}f(V_{t},Y_{t},t,\mu)\,dt+g(Y_{T},\mu)\Bigr]=\int_{U\times\mathbb{R}^{d}\times[0,T]}f(v,y,t,\mu)\Pi(dvdydt)\,dt+\int_{\mathbb{R}^{d}}g(y,\mu)\sigma_{T}(dy).$
@@
$\int_{U\times\mathbb{R}^{d}\times[0,T]}f(v,y,t,\mu)\,\Pi(dvdydt)\geq\int_{0}^{T}\int_{\mathbb{R}^{d}\times U}f(v(y,t),y,t,\mu)\sigma_{t}(dy)\,dt.$
According to the assertion (iii) of Theorem 2.1, the pair $(u,\mu)$ is optimal that is the following inequality
$\int_{0}^{T}\int_{\mathbb{R}^{d}}f(v(x,t),x,t,\mu)\,\sigma_{t}(dx)\,dt+\int_{\mathbb{R}^{d}}g(x,\mu)\sigma_{T}(dx)\geq\\ \int_{0}^{T}\int_{\mathbb{R}^{d}}f(u(x,t),x,t,\mu)\,\mu_{t}(dx)\,dt+\int_{\mathbb{R}^{d}}g(x,\mu)\mu_{T}(dx)$
holds. In terms of the processes $X_{t}$, $u(X_{t},t)$, $Y_{t}$ and $V_{t}$ the last inequality can be rewritten in the following way
$\mathbb{E}\Bigl[\int_{0}^{T}f(V_{t},Y_{t},t,\mu)\,dt+g(Y_{T},\mu)\Bigr]\geq\mathbb{E}\Bigl[\int_{0}^{T}f(u(X_{t},t),X_{t},t,\mu)\,dt+g(X_{T},\mu)\Bigr].$
Thus the statement (iii) of Corollary 2.1 is proved.
@@ Acknowledgements.
The authors are grateful to Prof. Vladimir I. Bogachev, Prof. Yurii V. Averboukh and Tikhon I. Krasovitskii for fruitful discussions and valuable remarks. We thank to the anonymous referee for thorough reading and useful corrections and suggestions.